OpenStudy (anonymous):

Can anyone help to solve the defferential equation. x^2dy/dx=9y

OpenStudy (anonymous):

u can solve it simply be seperating the variables and then integrating as follows (dy/y)=(9/(x^2)) dx on integrating we have ln y -ln c =-9/x (c being integrating constant) or ln(y/c) =-9/x or y/c =e^(-9/x) or y=c*e^(-9/x)